OpenStudy (anonymous):

Use implicit differentiation to find dy/dx x^3+3x^2y+y^3=8

OpenStudy (zehanz):

dy/dx=y' (easier to read, imo). Then: 3x²+6xy+3x²y'+3y²y'=0. Now try to isolate y' (solve for y').

OpenStudy (anonymous):

3x^2y'+3y^2y'=-3x^2-6xy ? then add like terms to 6x^2y'=-3x^2-6xy then divide by 6x^2 which equals \[\frac{ -3x^2-6xy }{ 6x^2 }\]

OpenStudy (zehanz):

I get: y'(3x²+3y²)=-3x²-6xy. So \(y'=\dfrac{-3x^2-6xy}{3x^2+3y^2}=-\dfrac{3x(x+2y)}{3(x^2+y^2)}=-\dfrac{x(x+2y)}{x^2+y^2}\)

OpenStudy (zehanz):

Thee are no like terms...

OpenStudy (anonymous):

ok I got it thanks

OpenStudy (zehanz):

YW!